# InSpire: Vision-Language-Action Models with Intrinsic Spatial Reasoning

### v2

@@ Appendix A Related Work
Learning Spatial Reasoning. Recently, a substantial focus has been placed on utilizing the spatial reasoning capabilities of VLMs or large language models (LLMs) to improve multimodal learning performance [25, 7, 18, 38, 12, 22, 17]. SpatialVLM [8] leverages 2D VLMs to understand spatial relationships and metric distances, effectively tackling the spatial relationship problem without an explicit 3D representation or scene graph. SpatialRGPT [18] advances VLMs’ spatial understanding by learning regional representation from 3D scene graphs and integrating depth information into the visual encoder. Our work is most aligned with SpatialVLA [47], which explores spatial representations for VLAs, by injecting 3D information into the input observations and representing spatial robot movement actions with adaptive discretized action grids. However, unlike SpatialVLA, which predicts actions based on encoded 3D positions and action grids, our proposed approach employs a visual question answering task as a bridging “language”, enabling simultaneous spatial reasoning and action prediction. This design endows VLAs with spatial reasoning capabilities without the use of extra data or interaction with other large models.
Appendix B Object Position Determination in Real-world Environments
To facilitate model training, ground-truth spatial relationships within visual observations are calculated based on the locations of the robot’s end-effector (or gripper) and task-specific objects. In simulation environments, object locations and gripper locations are readily available. In real-world environments, the end-effector’s position at the instant of gripper closure/opening is recorded and used as a proxy for the object’s position, as illustrated in Fig. 7.
Figure 7: Illustration of automatic object position determination in real-world environments. The end-effector’s position at the instant of gripper closure/opening is recorded and used as a proxy for the object’s position. The language instruction of the task is “put the banana in the plate”.
Appendix C Experimental Setup
C.1 Baseline Models
Simulation Experiments. Our proposed InSpire approach can be employed to improve existing autoregressive VLAs in a plug-and-play manner. We conduct simulation experiments on top of two state-of-the-art VLAs in the LIBERO environment, including miniVLA-VQ[3] and $\pi_{0}$-FAST[32].
  • MiniVLA-VQ [^4] https://ai.stanford.edu/blog/miniVLA: a 1B-parameter VLA that leverages a Qwen 2.5[40] 0.5B backbone while retaining the same ViT as OpenVLA [15] for visual encoding. The miniVLA-VQ model improves the vanilla miniVLA by integrating a vector quantization-based action chunking strategy to the policy learning process, achieving superior performance and computational efficiency compared with the strong competitor OpenVLA in the LIBERO[21] benchmark.
  • $\pi_{0}$-FAST [^5] https://www.physicalintelligence.company/research/fast: a 3B-parameter VLA trained via a compression-based tokenization approach that converts arbitrary robot action sequences into dense, discrete tokens for autoregressive VLA training. $\pi_{0}$-FAST changes the diffusion policy-based action prediction mode of $\pi_{0}$ [4] to autoregressive action prediction and accelerates training efficiency of $\pi_{0}$ considerably. By pretraining on a large and diverse dataset of robotic manipulation behaviors, both $\pi_{0}$-FAST and $\pi_{0}$ have demonstrated leading performance in a variety of real-world tasks.
Real-world Experiments. Given the challenges in collecting a large corpus of real-world behavioral demonstrations, it is difficult to train a model that displays generalization and robustness on varied real-world tasks, whether starting from scratch or finetuning a pretrained model lacking exposure to substantial real-world manipulation data. Therefore, we conduct real-world experiments using $\pi_{0}$-FAST, the current best-performing open-source VLA.
C.2 Evaluation Tasks
We perform simulation experiments using the LIBERO [21] benchmark [^6] https://LIBERO-project.github.io/datasets. Specifically, we utilize manipulation tasks from the five datasets—LIBERO-90, LIBERO-Spatial, LIBERO-Object, LIBERO-Goal and LIBERO-Long/10, which feature a diverse range of objects, scene layouts, and language instructions. LIBERO-Spatial, LIBERO-Object, LIBERO-Goal and LIBERO-Long/10 are respectively designed to scrutinize the controlled transfer of knowledge about spatial information, objects, task goals and planning abilities, and all have 10 tasks.
Table 4: Details of seen and unseen tasks used in our simulation and real-world environments.
Env. Seen Unseen
Simu. 90 LIBERO-90 tasks  10 LIBERO-Spatial tasks
 10 LIBERO-Object tasks
 10 LIBERO-Goal tasks
  10 LIBERO-Long tasks
Real. pick up the ball put the orange/yellow cup on the plate
pick up the plug put the ball on the book
move the can near the apple/watermelon stack the green lego on the blue lego
pull out the bottom/middle/top plate from the cabinet put the banana on the towel
push the bottom/middle/top plate into the cabinet pick up the orange
put the banana in the plate
put the blue/green/red cup in the plate
put the cookies on the towel
stack the green cube on the blue cube
stack the left/middle/right bowl on the left/middle/right bowl
Figure 8: We conduct 10 trials for each real-world task and randomize the configurations and orientations of objects for each trial. The language command for this task is “pick up the plug”.
We design 10 real-world manipulation tasks with diverse objects and scenes. For each task, we collect 10 demonstrations for training. In addition, we modify the language instructions and corresponding objects of the 10 original tasks to construct 5 unseen tasks. The details of these tasks are presented in Table 4. In alignment with the simulation experiments, the model is finetuned on seen tasks, and evaluated on both seen and unseen tasks. During evaluation/inference, we conduct 10 trials for each real-world task. We randomize the configurations as well as orientations of task-specific objects for each trial, as illustrated in Fig. 8.
C.3 Training Hyperparameters
In our simulation experiments, the baseline model miniVLA-VQ is pretrained from scratch while $\pi_{0}$-FAST is finetuned on tasks from the LIBERO-90 dataset, the hyperparameters for training miniVLA-VQ and $\pi_{0}$-FAST are listed in Table 5 and Table 6, respectively. In our real-world experiments, the hyperparameters for training $\pi_{0}$-FAST are identical to those used in the simulation experiments. To ensure fair comparisons, our InSpire method uses the same hyperparameters as those employed by the two baseline VLAs. The hyperparameters for finetuning SpatialVLA [47] on tasks from LIBERO-Spatial, LIBERO-Object, LIBERO-Goal and LIBERO-Long is listed in Table 7.
Table 5: Hyperparameters for training miniVLA-VQ [3] on LIBERO-90 tasks.
Hyperparametes Setting
batch size 128
optimizer adamW
learning rate 1e-5
weight decay 0
training step 50000
GPUs 4$\times$×A800
Table 6: Hyperparameters for finetuning $\pi_{0}$-FAST [32] on LIBERO-90 or real-world tasks.
Hyperparametes Setting
batch size 128
optimizer adamW
learning rate 2.5e-5
lr schedule cosine decay
warmup step 1000
weight decay 1e-10
training step 30000
action chunk 5
GPUs 4$\times$×A800
Table 7: Hyperparameters for finetuning SpatialVLA [47] on tasks from LIBERO-Spatial, LIBERO-Object, LIBERO-Goal and LIBERO-Long.
Hyperparametes Setting
batch size 128
optimizer adamW
learning rate 2.5e-4
lr schedule linear
weight decay 0
training step 30000
lora rank 32
action chunk 4
GPUs 4$\times$×A800
Figure 9: Performance at different training steps.
Appendix D Additional Results
Figure 10: Attention maps of the baseline model miniVLA-VQ [3] integrated w/ or w/o our InspireVLA approach on various manipulation tasks.
D.1 Qualitative Analysis.
Fig. 10 presents the attention maps of the baseline model miniVLA-VQ integrated w/ or w/o our InspireVLA approach on various manipulation tasks. As shown in the figure, the baseline VLA tends to predict actions based on spurious or background features in visual observations. InSpire effectively enhance the baseline VLA by redirecting its attention to task-relevant regions.
D.2 Performance at Different Training Steps
Fig. 9 presents the performance of the baseline model and our InSpire scheme on seen and unseen tasks across different training epochs. As illustrated, our InSpire consistently enhances the success rates of the baseline across all training epochs and exhibits strong training stability.
D.3 Detailed Ablation Results
We report the success rates (%) of our proposed InSpire approach equipped with different VQA formulations and VQA insertion positions in Tables 8 and 9, respectively. The baseline model in the table is miniVLA-VQ. We conduct 10 trials per task. As can be observed, InSpire consistently improves the baseline on seen tasks from the LIBER-90 dataset as well as unseen tasks from the datasets of LIBERO-Spatial, LIBERO-Object, LIBERO-Goal and LIBERO-Long.
Table 8: Success rates (%) of Inspire with different VQA formulations.
Model Seen Unseen
LIBERO-90 LIBERO-Spatial LIBERO-Object LIBERO-Goal LIBERO-Long Average
Baseline 83.8 0 0 5 9 3.5
1D Direction 90.8 23 27 12 10 18.0
3D Direction 90.0 9 28 21 10 16.9
Proximity 88.7 4 18 10 5 9.3
3D Location 88.9 2 20 12 6 10.1
Distance 82.8 1 4 13 6 6.0
Table 9: Success rates (%) of Inspire with various VQA insertion positions.
Model Seen Unseen
LIBERO-90 LIBERO-Spatial LIBERO-Object LIBERO-Goal LIBERO-Long Average
Baseline 83.8 0 0 5 9 3.5
VQA-First 90.8 23 27 12 10 18.0
Instruct-First 88.7 18 34 16 10 19.5

### v3

@@ IV-3 Learning Spatial Reasoning
Leveraging the spatial reasoning of large models is a recent focus in multimodal learning [25, 7, 21, 36, 12, 23, 19]. Methods range from using 2D VLMs for spatial understanding without explicit 3D data (SpatialVLM [8]) to integrating 3D scene graphs and depth information (SpatialRGPT [21]). Our work is most related to SpatialVLA [44], which injects 3D information and uses discretized action grids. However, instead of relying on encoded 3D positions, our approach employs a VQA task as a bridging “language” to enable simultaneous spatial reasoning and action prediction, endowing VLAs with this capability without extra data or models.
V Conclusion, Limitations and Future Work
This work proposes Intrinsic Spatial Reasoning (InSpire) to mitigate the adverse effects of spurious correlations on the action prediction of VLAs. By incorporating a spatial reasoning VQA task as the bridging “language” between visual observations and low-level actions, InSpire endows VLAs with spatial reasoning capabilities without employing extra data or interacting with other large models. Notably, InSpire can be used as a plugin to improve existing autoregressive VLAs. Comprehensive experiments demonstrate InSpire’s effectiveness and flexibility, achieving consistent improvements on two state-of-the-art VLAs on both seen and unseen tasks in simulation and real-world environments.
Although the results are encouraging, two key limitations remain in this work: 1) Although InSpire has potential for integration with any LLM-based VLAs, its effectiveness in enhancing diffusion policy-based VLAs, such as Octo [26] and $\pi_{0}$ [4], remains to be explored. Future work will focus on exploring strategies to facilitate seamless integration of InSpire into such models. 2) VLAs pretrained with diverse paradigms and datasets likely acquire varied spurious correlations; however, the characters of these differences across models are not yet understood. A deeper empirical study of these factors is a promising direction for bolstering the robustness of our approach in more complex scenarios.